OpenStudy (anonymous):

The traffic cone is 19 inches tall & has a radius of 5 inches. Find the surface area.

OpenStudy (marissalovescats):

Hi and welcome to OpenStudy! :) First off we're going to want to know the surface area of a cone and that is:\[SA = \Pi r(r+\sqrt{h^2+r^2)}\]

OpenStudy (marissalovescats):

Where r is the radius and h is the height So we plug in the numbers where they go and solve!

OpenStudy (marissalovescats):

\[SA = \Pi (5)(5+\sqrt{(16^2+5^2)}\]

OpenStudy (marissalovescats):

16^2 we know is 256 and 5^2 is 25 so :\[SA = \Pi (5)(5+\sqrt{(256+25)}\]

OpenStudy (marissalovescats):

256 +25 is 281 and the square root of 281 is about 16.76 and 16.76+5 is 21.76 so now we have this: \[SA = \Pi (5)(21.76)\]

OpenStudy (marissalovescats):

So our surface area is about: \[SA = 108.82 \Pi \] Which is also 341.85 in^2